OpenStudy (anonymous):

if the sum of 2 numbers is 70. Their ratio is 3:7. what is the larger number?

ganeshie8 (ganeshie8):

larger number = \(\frac{7}{3+7}\times 70\)

OpenStudy (whpalmer4):

\[a + b =70\]\[7a = 3b\] solve by substitution. or, what @ganeshie8 did was say that the smallest combination of two numbers in that ratio is 3 and 7, and added together they equal 10. then scale it up by a factor of 7 so that together they equal 70.

OpenStudy (whpalmer4):

\[\frac{3}{7}*\frac{7}{7}= \](don't simplify, just look at the result)